# Semialgebraic Neural Networks: From roots to representations

### v1

@@ G.3 Semialgebraic transformer
In this section, we demonstrate how a variation of the transformer architecture Vaswani et al. [2023] is expressible as a SANN.
@@
Figure 6: $\mathrm{saxp}$ (black) is a semialgebraic approximation to $\exp$ (red). It is simple, reasonably accurate near 00, and has the correct asymptotic behavior as $x\to\pm\infty$.
$\mathrm{saxp}$ is a good approximation of $\exp$ near 00, but $\mathrm{saxp}=o(\exp)$ as $x\to\infty$. In fact this is true for every semialgebraic function since every semialgebraic function is bounded above by a polynomial [Bochnak and Efroymson, 1980]. This is not a concern for our transformers since layer normalization will keep the input to $\mathrm{saxp}$ near 00.
@@
Recall that every function defined piecewise on a semialgebraic decomposition of $\mathbb{R}^{m}$ using only addition, subtraction, multiplication, division, and extraction of roots is a semialgebraic functions. From this, it is clear that $\mathrm{MAtt}$, $\mathrm{MLP}$, $\mathrm{LayerNorm}$ in equations (99), (100), and (101) are semialgebraic. Furthermore, they are all continuous, so they are bounded when their domain is restricted to be compact, such as after a tokenization step. Finally, note that the composition of semialgebraic functions is itself a semialgebraic function.

### v2

@@ G.3 Semialgebraic transformer
In this section, we demonstrate how a variation of the transformer architecture (Vaswani et al., 2023) is expressible as a SANN. Our presentation of the transformer architecture is inspired by Furuya et al. (2024).
@@
Figure 6: $\mathrm{saxp}$ (black) is a semialgebraic approximation to $\exp$ (red). It is simple, accurate near 00, and has the correct asymptotic behavior as $x\to\pm\infty$.
$\mathrm{saxp}$ is a good approximation of $\exp$ near 00, but $\mathrm{saxp}=o(\exp)$ as $x\to\infty$. In fact this is true for every semialgebraic function since every semialgebraic function is bounded above by a polynomial (Bochnak & Efroymson, 1980). This is not a concern for our transformers since layer normalization will keep the input to $\mathrm{saxp}$ near 00.
@@
Recall that every function defined piecewise on a semialgebraic decomposition of $\mathbb{R}^{m}$ using only addition, subtraction, multiplication, division, and extraction of roots is a semialgebraic function. From this, it is clear that $\mathrm{MAtt}$, $\mathrm{MLP}$, and $\mathrm{LayerNorm}$ in equations (98), (99), and (100) are semialgebraic. Furthermore, they are all continuous, so they are bounded when their domain is restricted to be compact, such as after a tokenization step. Finally, note that the composition of semialgebraic functions is itself a semialgebraic function.